OpenStudy (anonymous):

what is the vertex of the function f(x)=2x^2+4x-16? Write the anser in the form (x,y)

sam (.sam.):

Complete the square

sam (.sam.):

or use -b/2a

OpenStudy (anonymous):

i don't get it

sam (.sam.):

2x^2+4x-16 2(x^2+2x)-16 2(x+1)^2-1(2)-16 2(x+1)^2-18 vertex at (-1,-18)

OpenStudy (anonymous):

thank you....i'm trying to get what you did. You added 1. Why?

OpenStudy (anonymous):

ok...i see. i had

sam (.sam.):

Its completing the square , did you learn it?

OpenStudy (anonymous):

the teacher went over it but i just couldn't get it!! I think I have it a little better now. at least i hope before my final.

sam (.sam.):

The technique goes by ax^2+bx+c=0 The coefficient of x^2 must be 1, so factor 'a' out a(x^2+bx/a)+c=0 Then here's the technique, divide bx/a by 2,then set the x^2 becomes x only,eliminate the x from b/a, then square the whole bracket you ended up with a(x+b/2a)^2+c=0 Still haven't finished, you need to bring the b/2a OUT the brackets and square it, and when you bring it out, it is always negative, be sure that the 'a' will be multiplying the one you bought it out a(x+b/2a)^2-(b/2a)^2(a)+c=0